I am trying to create an algebric expression from this worded problem
One angle is 6 degrees less than twice a second angle. If the two angles are complementary, what are their measures?

Answer by mananth(16949)

let the angles be x & y
x+y =90
x= 2y-6
substitute x in the first equation
2y-6+y = 90
3y=96
/3
y=32
now you know what x is